Not sure how to do this f(x) inverse function:
m(x)=1/x+3
Please explain,
Thank You

Answer by lwsshak3(11628)

Not sure how to do this f(x) inverse function:
m(x)=1/x+3
y=1/x+3
interchange x and y, then solve for y
x=1/y+3
x-3=1/y
y=1/(x-3)
y^-1=1/(x-3)